### v1

# Think2SQL: Reinforce LLM Reasoning Capabilities for Text2SQL

Abstract
Large Language Models (LLMs) have shown impressive capabilities in transforming natural language questions about relational databases into SQL queries. Despite recent improvements, small LLMs struggle to handle questions involving multiple tables and complex SQL patterns under a Zero-Shot Learning (ZSL) setting. Supervised Fine-Tuning (SFT) partially compensate the knowledge deficits in pretrained models but falls short while dealing with queries involving multi-hop reasoning. To bridge this gap, different LLM training strategies to reinforce reasoning capabilities have been proposed, ranging from leveraging a thinking process within ZSL, including reasoning traces in SFT, or adopt Reinforcement Learning (RL) strategies. However, the influence of reasoning on Text2SQL performance is still largely unexplored.
This paper investigates to what extent LLM reasoning capabilities influence their Text2SQL performance on four benchmark datasets. To this end, it considers the following LLM settings: (1) ZSL, including general-purpose reasoning or not; (2) SFT, with and without task-specific reasoning traces; (3) RL, leveraging execution accuracy as primary reward function; (4) SFT+RL, i.e, a two-stage approach that combines SFT and RL.
The results show that general-purpose reasoning under ZSL proves to be ineffective in tackling complex Text2SQL cases. Small LLMs benefit from SFT with reasoning much more than larger ones, bridging the gap of their (weaker) model pretraining. RL is generally beneficial across all tested models and datasets, particularly when SQL queries involve multi-hop reasoning and multiple tables. Small LLMs with SFT+RL excel on most complex datasets thanks to a strategic balance between generality of the reasoning process and optimization of the execution accuracy. Thanks to RL, the 7B Qwen-Coder-2.5 model performs on par with 100+ Billion ones on the Bird dataset.
1 Introduction
The ever-increasing volume of data stored in relational databases and the impressive diffusion of Large Language Models (LLMs) have jointly paved the way for new accessible ways to query multi-table databases. The Text2SQL task involves converting natural language questions about relational tables into executable SQL queries [12]. Thanks to Text2SQL models, end-users who are not proficient in SQL coding can simply access relational data by using LLMs as a proxy. Tackling Text2SQL is particularly challenging as not only involves expressing first- or second-order logic conditions in SQL but also reasoning about the underlying question’s meaning and its relation to the database schema [13].
Neural network-based solutions to Text2SQL have evolved from classical sequence-to-sequence and graph networks (e.g., [54, 2]) to Transformer-based architectures [1] and, more recently, to LLM-based solutions [27]. Thanks to the advanced language understanding capabilities of their pretrained models, LLMs have remarkably boosted Text2SQL performance, particularly on multi-table datasets [26, 56]. However, LLMs’ performance under Zero-Shot Learning (ZSL) significantly varies depending on the number of model parameters [5]. While small LLMs (i.e., models with 3-8 Billions of parameters) suffer from limited language understanding and reasoning capabilities, larger ones are typically trained on multi-domain data thus lacking the adequate level of specialization to be competitive on domain-specific data [6].
To overcome the limitations of ZSL, Supervised Fine-Tuning (SFT) is among the mostly used LLM adaptation strategies [52]. It entails specializing the language model parameters for a given downstream task, such as Text2SQL. Since SFT requires task-specific data (e.g., pairs of natural language questions and the corresponding SQL queries), the curation of an annotated SQL-centric corpus [25] is critical. Moreover, even when training data and resources are appropriate, small LLMs typically show limited generality and reasoning capabilities, especially while coping with complex database schema and SQL patterns [36].
Reinforcement Learning (RL) techniques have recently proved to be the most effective in improving LLM reasoning capabilities [10]. Although this LLM training strategy have led to state-of-the-art results in several downstream tasks, such as mathematical reasoning and Python/Java coding, its influence on Text2SQL performance is still largely unexplored.
In this paper, we thoroughly analyze the influence of LLM reasoning capabilities on Text2SQL performance. To achieve this goal, we evaluate both pretrained LLMs with thinking capabilities and LLMs specialized for reasoning on Text2SQL under the following settings:
Zero-Shot Learning (ZSL) with general-purpose reasoning. We consider pretrained LLMs (e.g., [41, 21]) that already incorporate the reasoning steps in their pretrained model, but are not specifically suited to the Text2SQL task.
Supervised Fine-Tuning (SFT) with reasoning. We fine-tune small LLMs for the Text2SQL task. We prepare a task-specific dataset covering SQL patterns with varying levels of complexity. SFT examples are enriched with reasoning traces to make the problem solving step explicit to the LLM during training.
Reinforcement Learning (RL). We tailor RL to the Text2SQL task. The LLM repetitively performs actions consisting of shortlisting the best SQL query to solve the input question among a predefined set of candidates. We adopt Group-Relative Policy Optimization (GRPO) [41] and reward the LLM answers by using the established execution accuracy metric [56].
Supervised Fine-Tuning and Reinforcement Learning (SFT+RL). We employ a two-step approach combining SFT with RL. The idea behind it is to specialize the model on problem solving using RL while keeping the generality of reasoning models [10].
Our evaluation aims to address the following Research Questions (RQs):
  • RQ1)
    Is reasoning beneficial for Text2SQL performance under different LLM training settings?
  • RQ2)
    Which is the most appropriate strategy to train LLMs to reason about Text2SQL?
  • RQ3)
    Which is the best trade-off between model generalization and specialization?
To answer RQ1, we compare the results of LLMs under the ZSL setting with and without reasoning as well as the performance of LLMs under the SFT setting with and without reasoning traces. The goal is to explore the influence of reasoning on the performance of LLMs with different numbers of parameters, pretraining strategies, and across testing datasets with different characteristics.
To answer RQ2, we compare the performance of LLMs under (1) ZSL with reasoning vs. (2) SFT with reasoning vs. (3) RL vs. (4) SFT+RL. The goal is to compare different strategies to incorporate reasoning capabilities in LLM training for Text2SQL, with particular attention to the model performance achieved on complex SQL patterns.
To answer RQ3, we analyze the LLMs’ performance across diverse datasets, ranging from general-purpose to domain-specific content. The purpose is to clarify whether reasoning is beneficial to achieve model generality across different datasets and domains while preserving the overall accuracy of the SQL query generator.
The rest of the paper is organized as follows. Section 2 introduces the preliminary notions and the Text2SQL problem formulation. Section 3 describes the methodology used to assess the influence of reasoning on Text2SQL performance. Section 4 describes the experimental settings and the main results. Finally, Section 5 draws conclusions and discusses limitations and future extensions of the present work.
2 Preliminaries
In this section, we introduce the notation and fundamental formulations used throughout this work. Let a sequence of discrete tokens be represented as ${{\boldsymbol{z}}}=(z_{1},z_{2},\ldots,z_{T}),$ where each $z_{t}\in\mathcal{V}$, and $\mathcal{V}$ is a finite vocabulary set with cardinality $V$. We consider a large language model (LLM) parameterized by ${{\boldsymbol{\theta}}}$, formalized as a probabilistic autoregressive model $\pi_{{{\boldsymbol{\theta}}}}$, instantiated as a decoder-only transformer architecture [38]. For a given sequence ${{\boldsymbol{z}}}$, the model defines a factorized distribution over the sequence space:
$\pi_{{{\boldsymbol{\theta}}}}({{\boldsymbol{z}}})=\prod_{t=1}^{T}\pi_{{{\boldsymbol{\theta}}}}(z_{t}\mid{{\boldsymbol{z}}}_{<t}),$ (1)
where ${{\boldsymbol{z}}}_{<t}=(z_{1},\ldots,z_{t-1})$ is the left-truncated context of length $t-1$. Each conditional probability $\pi_{{{\boldsymbol{\theta}}}}(z_{t}\mid{{\boldsymbol{z}}}_{<t})$ is computed via a series of masked multi-head self-attention layers. Causal masking ensures that attention weights for token $z_{t}$ are computed only over ${{\boldsymbol{z}}}_{<t}$, preserving the autoregressive property. All model parameters, including the token embeddings, attention weights, feedforward weights, normalization scales, and the output projection matrix, are collected in ${{\boldsymbol{\theta}}}$.
2.1 Fine-Tuning Language Models
Supervised Fine-Tuning (SFT) adapts a pretrained language model $\pi_{{{\boldsymbol{\theta}}}}$ to a distribution $\mathcal{P}$ of sequences that reflect desired linguistic or task-specific behavior. Let ${{\boldsymbol{z}}}=(z_{1},z_{2},\ldots,z_{T})$ denote a token sequence drawn from ${{\boldsymbol{z}}}\sim\mathcal{P}$. The SFT objective maximizes the likelihood of sequences under $\pi_{{{\boldsymbol{\theta}}}}$, which corresponds to minimizing the expected negative log-likelihood:
$\mathcal{L}_{\text{full}}({{\boldsymbol{\theta}}})=-\mathbb{E}_{{{\boldsymbol{z}}}\sim\mathcal{P}}\left[\sum_{t=1}^{T}\log\pi_{{{\boldsymbol{\theta}}}}(z_{t}\mid{{\boldsymbol{z}}}_{<t})\right].$ (2)
For tasks with an explicit decomposition into an input segment and a target segment—such as QA, summarization, or assistant-style dialogue—the data distribution consists of pairs $({{\boldsymbol{x}}},{{\boldsymbol{y}}})\sim\mathcal{P}$, where ${{\boldsymbol{x}}}=(x_{1},\ldots,x_{n})$ is the conditioning prompt and ${{\boldsymbol{y}}}=(y_{1},\ldots,y_{m})$ is the supervised output. In such settings, the model conditions on ${{\boldsymbol{x}}}$ and predicts the continuation ${{\boldsymbol{y}}}$, with the loss computed only over the target tokens:
$\mathcal{L}_{\text{cond}}({{\boldsymbol{\theta}}})=-\mathbb{E}_{({{\boldsymbol{x}}},{{\boldsymbol{y}}})\sim\mathcal{P}}\left[\sum_{t=1}^{m}\log\pi_{{{\boldsymbol{\theta}}}}(y_{t}\mid{{\boldsymbol{x}}}\|{{\boldsymbol{y}}}_{<t})\right]\,,$ (3)
[…]
2.2 Reinforcement Learning for Language Models
Reinforcement Learning from Human Feedback (RLHF) typically relies on policy optimization algorithms to fine-tune a language model $\pi_{{{\boldsymbol{\theta}}}}$ toward reward-aligned behavior. One of the standard approaches is Proximal Policy Optimization (PPO) [40], which constrains policy updates through a clipped surrogate objective and value-based advantage estimation. However, PPO necessitates learning and maintaining an auxiliary value function, which introduces instability and potential reward misestimation.
Group-Relative Policy Optimization (GRPO) [41] offers a value-free alternative by computing normalized group-level advantages based directly on realized rewards. Let ${{\boldsymbol{x}}}\sim\mathcal{X}$ denote a prompt drawn from a distribution over conditioning inputs, and let $\{{{\boldsymbol{y}}}_{i}\}_{i=1}^{G}\sim\pi_{{{\boldsymbol{\theta}}}_{\text{old}}}(\cdot\mid{{\boldsymbol{x}}})$ be $G$ response sequences generated by the frozen reference policy $\pi_{{{\boldsymbol{\theta}}}_{\text{old}}}$. Each response ${{\boldsymbol{y}}}_{i}=(y_{i,1},\ldots,y_{i,T_{i}})$ is assigned a scalar reward $R_{i}\in\mathbb{R}$ computed via a reward model.
The group-relative advantage $A_{i}$ for the $i$-th response is defined by normalizing the reward distribution over the group:
$A_{i}=\frac{R_{i}-\mathbb{E}[R_{j}]}{\sqrt{\mathbb{V}[R_{j}]}},\qquad j\in\{1,\ldots,G\},$ (4)
where $\mathbb{E}[R_{j}]$ and $\mathbb{V}[R_{j}]$ are the mean and variance of the rewards for the group of responses, respectively. For each token position $t$ in response ${{\boldsymbol{y}}}_{i}$, define the state as ${{\boldsymbol{s}}}_{i,t}={{\boldsymbol{x}}}\|{{\boldsymbol{y}}}_{i,<t}$, and the token-level probability ratio as
$p_{i,t}({{\boldsymbol{\theta}}})=\frac{\pi_{{{\boldsymbol{\theta}}}}(y_{i,t}\mid{{\boldsymbol{s}}}_{i,t})}{\pi_{{{\boldsymbol{\theta}}}_{\text{old}}}(y_{i,t}\mid{{\boldsymbol{s}}}_{i,t})}.$
The GRPO training objective minimizes a clipped surrogate loss penalized by the KL divergence from the reference policy:
$\mathcal{L}_{\text{GRPO}}({{\boldsymbol{\theta}}})=\mathbb{E}\left[\frac{1}{G}\sum_{i=1}^{G}\frac{1}{T_{i}}\sum_{t=1}^{T_{i}}\min\left(p_{i,t}({{\boldsymbol{\theta}}})A_{i},\;\text{clip}\big{(}p_{i,t}({{\boldsymbol{\theta}}}),1-\epsilon,1+\epsilon\big{)}A_{i}\right)-\beta\textsc{kl}\left[\pi_{{{\boldsymbol{\theta}}}}\;\|\;\pi_{{{\boldsymbol{\theta}}}_{\text{ref}}}\right]\right],$ (5)
where the expectation is taken over the prompt distribution ${{\boldsymbol{x}}}\in\mathcal{X}$, and the responses $\{{{\boldsymbol{y}}}_{i}\}_{i=1}^{G}$ generated by the frozen policy $\pi_{{{\boldsymbol{\theta}}}_{\text{old}}}$. Additionally, $\epsilon$ is set to be the clipping parameter and $\beta$ controls the Kullback-Leibler divergence (KL) regularization by penalizing models that deviate from the reference policy $\pi_{{{\boldsymbol{\theta}}}_{\text{ref}}}$ (which is typically the initial pretrained model).
2.2.1 Rule-based Reward Modeling
Reward modeling is central to reinforcement learning with language models, as it defines the optimization signal guiding the policy $\pi_{{{\boldsymbol{\theta}}}}$. Learned neural reward models are commonly employed to approximate human preferences or task-specific goals. However, they often suffer from distributional mismatch, reward hacking, and spurious correlations [17, 53, 11]. These effects arise when the model exploits imperfections in the reward predictor, leading to high-reward outputs that do not correspond to true task success.
An alternative is to design rule-based reward models, which define deterministic mappings from model outputs to scalar reward values via explicit criteria. In the context of coding, for instance, a reward function $R:{{\boldsymbol{y}}}\mapsto[0,1]$ can be constructed by executing the generated code ${{\boldsymbol{y}}}$ against a test suite and returning the fraction of passed unit tests. Such rule-based models directly encode correctness and task satisfaction, avoiding pathologies introduced by learned approximators.
Formally, let ${{\boldsymbol{x}}}\sim\mathcal{X}$ denote the input (e.g., a natural language instruction), and ${{\boldsymbol{y}}}\sim\pi_{\theta_{\text{old}}}(\cdot\mid{{\boldsymbol{x}}})$ a candidate response. The reward function $R({{\boldsymbol{x}}},{{\boldsymbol{y}}})\in\mathbb{R}$ is defined deterministically via evaluation procedures specified a priori. These functions are task-dependent and vary across application domains. The resulting reward is used to construct advantage estimates, as in GRPO.
A well-known limitation of rule-based reward models is the sparsity of the reward signal. In many structured tasks, the reward $R({{\boldsymbol{x}}},{{\boldsymbol{y}}})$ may remain zero across most model outputs and attain nonzero values only when the generation exactly satisfies task constraints. This sparsity complicates credit assignment during training and may impair exploration in RL-based optimization. Techniques such as reward shaping, curriculum learning, or relaxed matching criteria are sometimes introduced to mitigate this issue [31, 44, 17, 30]. Nonetheless, provided that the policy starts from a sufficiently strong pretrained model, this approach has been successfully adopted in multiple recent frameworks across general and specialized RLHF pipelines [10, 46, 55], and has been particularly effective in settings where ground truth verification criteria exist, such as program synthesis [23, 18, 8].
2.3 Text2SQL
The Text2SQL task consists in mapping natural language question on databases to executable SQL code. Let ${{\boldsymbol{x}}}\in\mathcal{X}$ denote a natural language input (e.g., a user question), and let ${{\boldsymbol{y}}}\in\mathcal{Y}_{\text{sql}}$ denote a corresponding structured output in SQL syntax. The output space $\mathcal{Y}_{\text{sql}}$ comprises syntactically valid SQL codes consistent with a given database schema $\mathcal{S}$, which specifies the collection of relational tables, attributes, and their types. In addition to the schema, auxiliary context $\mathcal{M}$ may be provided. This includes task-specific metadata such as database descriptions, natural language annotations, examples of prior queries, or column-level summaries.
The schema is tokenized into a unified model-readable representation via a a deterministic transformation $\phi:\mathcal{S}\mapsto\mathcal{V}$, which converts the schema $\mathcal{S}$ into a token sequence $\phi(\mathcal{S})$ compatible with the model’s input vocabulary. For this study, we adopt the schema representation prompt commonly used in prior works for its proven effectiveness [4, 16].
The model $\pi_{{{\boldsymbol{\theta}}}}$ defines a conditional distribution over the SQL code ${{\boldsymbol{y}}}=(y_{1},\ldots,y_{T})$ given the input ${{\boldsymbol{x}}}$ and the schema context:
$\pi_{{{\boldsymbol{\theta}}}}({{\boldsymbol{y}}}\mid{{\boldsymbol{x}}},\phi(\mathcal{S}),\mathcal{M})=\prod_{t=1}^{T}\pi_{{{\boldsymbol{\theta}}}}(y_{t}\mid{{\boldsymbol{x}}}\|\phi(\mathcal{S})\|\mathcal{M}\|{{\boldsymbol{y}}}_{<t}).$ (6)
For each prompt ${{\boldsymbol{x}}}$, the target may be a set of logically equivalent parses $\mathcal{Y}^{*}({{\boldsymbol{x}}})\subseteq\mathcal{Y}_{\text{sql}}$, all yielding identical execution results. Learning may proceed by optimizing the marginal log-likelihood over this set or by selecting a canonical representative from $\mathcal{Y}^{*}({{\boldsymbol{x}}})$ during training [58, 50, 43]. In this study, the latter approach is adopted, and the model is trained to predict a single SQL code ${{\boldsymbol{y}}}^{*}\in\mathcal{Y}^{*}({{\boldsymbol{x}}})$ for each input ${{\boldsymbol{x}}}$.
To manage large schemas, modern systems restrict $\phi(\mathcal{S})$ to a localized substructure $\phi(\mathcal{S}_{{{\boldsymbol{x}}}})$, where $\mathcal{S}_{{{\boldsymbol{x}}}}\subseteq\mathcal{S}$ is retrieved via schema linking, lexical overlap, or learned attention [49, 37, 45, 7, 3]. To better isolate the reasoning process in Text2SQL and disentangle schema linking from SQL generation, we restrict $\mathcal{S}_{{{\boldsymbol{x}}}}$ during both training and inference to include only the tables and their complete schema that are directly relevant to the question ${{\boldsymbol{x}}}$.
3 Methodology
To evaluate the influence of reasoning on the Text2SQL task, we employed several training strategies. These strategies include supervised fine-tuning (SFT), reinforcement learning (RL), and a hybrid approach combining both. Each strategy was designed to assess the impact of reasoning traces on model performance.
Supervised Fine-Tuning (SFT). In the SFT approach, the model was trained on the curated dataset described in Section 3.1. The training objective was to minimize the cross-entropy loss between the predicted SQL code and the ground truth SQL code. Reasoning traces were included as additional input to guide the model in understanding the logical steps required to generate the correct SQL.
Reinforcement Learning (RL). For RL, we used execution accuracy as the primary reward signal. The model was fine-tuned using the GRPO algorithm, where the reward was computed based on the correctness of the generated SQL code’s execution results. To encourage the generation of reasoning traces, we also included a secondary reward signal based on syntatical checks. Details on the rewards are provided in Section 3.2.
Hybrid Approach. The hybrid approach combined SFT and RL. The model was first trained using SFT to leverage the labeled dataset and then fine-tuned with RL to further optimize execution accuracy and reasoning quality. This two-stage training process aimed to balance the benefits of supervised learning and reinforcement learning.
3.1 SFT Dataset Creation
This section describes the creation of a complex reasoning dataset tailored for the Text2SQL task. Text2SQL was chosen due to its practical significance, its prominence in recent advancements [19], and its familiarity to large language models (LLMs). LLMs have shown strong performance on established benchmarks [56, 26, 24]. Furthermore, SQL queries, unlike some other logical forms, can be executed and verified for correctness, making them particularly suitable for this study.
Data Collection. The initial phase of our methodology involved the acquisition of high-quality, human-annotated Text2SQL datasets. For this study, we selected the BIRD dataset [26], recognized for its extensive scope and diversity.
The BIRD training set comprises 9,428 data points derived from 69 heterogeneous databases spanning 37 professional domains, including blockchain, healthcare, education, and hockey. Each data point consists of a natural language (NL) question, a corresponding SQL code, and supplementary evidence. The evidence serves as additional context to resolve ambiguities in the schema or NL questions.
Data Quality and Complexity Curation. To ensure the reliability and robustness of the dataset, we implemented a rigorous two-step curation process [29]. First, we filtered out erroneous SQL queries and removed duplicate entries, resulting in the exclusion of 421 instances. Second, we categorized the remaining SQL queries based on their complexity, defined by the number of SQL constructs, into three tiers: low ($[0,7)$), medium ($[7,10)$), and high ($[10,+\infty)$). This stratification yielded a dataset distribution of 7,022 simple-complexity queries (77%), 1,549 medium-complexity queries (17%), and 492 challenging queries (6%).
Prompt Synthetic data annotation Answer the following question with the SQL code. Use the piece of evidence and base your answer on the database schema. Given the question, the evidence and the database schema, return in the answer tags only the SQL script that addresses the question. Question: <question> Evidence: <evidence> Database Schema: <schema>
Figure 1: Prompt used for the synthetic data annotation. <question>, <evidence>, and <schema> are placeholders for the actual question, evidence, and database schema, respectively. The model is expected to generate a SQL code snippet that answers the question based on the provided evidence and schema.
Synthetic Data Annotation. To enhance the dataset with reasoning traces, we utilized the DeepSeek-R1 model [10] along with its system prompt. The prompt used for synthetic annotation is detailed in Figure 1. The hyperparameters optimized for reasoning tasks were set to a temperature of 0.7 and a top-p of 0.95, following established best practices in the field [48, 32]. This configuration ensured high-quality annotations that explicitly outline reasoning steps, enriching the dataset and enabling the model to better comprehend the logical processes underlying SQL generation.
The final annotated dataset consists of 1,142 instances, distributed as follows: 684 simple queries, 265 medium-complexity queries, and 193 challenging queries. The 75th percentile of reasoning token counts is 509 for simple queries, 861 for medium queries, and 869 for challenging queries. The dataset will be made available on Hugging Face [^1] To appear soon.
3.2 Rewards for Reinforcement Learning
In reinforcement learning, reward signals are crucial for guiding the model’s learning process [17, 53, 11]. Execution accuracy, the primary reward for Text2SQL, measures the correctness of generated SQL by comparing it to the ground truth. However, its binary nature poses challenges for RL optimization, especially for smaller LLMs, as rewards often remain zero unless the SQL is exactly correct. To encourage the model’s reasoning process, we introduce an auxiliary reward signal that evaluates the correctness of the reasoning trace based on the appropriate use of reasoning tags [10]. Additionally, to mitigate reward hacking, a penalty is applied when reasoning tokens are redundantly or excessively repeated within the reasoning trace.
Execution Accuracy (EX). Execution accuracy [56, 26] evaluates the correctness of the generated SQL by comparing its execution results against those of the ground truth SQL on the target database. This metric provides a binary reward, assigning a full score only when the two execution results match exactly, row by row. While straightforward and reliable, execution accuracy has a notable limitation: it does not account for partially correct results, such as queries that produce subsets or supersets of the expected results, which can hinder the learning process in reinforcement learning scenarios.
Format Reward. The format reward [10] incentivizes the model to adhere to a predefined output structure, such as the use of <think> and <answer> tags. Outputs that conform to this structure receive a reward boost, promoting clarity and consistency. This is a sparse reward that activates only when both the opening and closing tags for reasoning and answers are correctly positioned. The reward value is 1 if the tags are correctly formatted; otherwise, it is 0.
Tag Count Reward. To address reward hacking, where reasoning traces include unnecessary or excessive tags, we introduce a tag count reward. This reward penalizes the model for generating reasoning traces with redundant tags. The reward is 1 if the number of tags is exactly two (one opening and one closing tag); otherwise, it is 0.
The final reward signal is computed as a weighted sum of the individual rewards. The weights were carefully chosen to ensure a balanced contribution from each reward component while maintaining a total score of 1. This design prevents training instability caused by excessively high rewards. The final reward is computed as follows:
$R=0.85\cdot R_{\text{EX}}+0.10\cdot R_{\text{Format}}+0.05\cdot R_{\text{Tag Count}},\qquad R\in[0,1]$ (7)
Table 1: Performance comparison of open-source and proprietary models on the Bird Dev dataset for the Text2SQL task. All models were evaluated with a temperature setting of 0.7 and a top_p value of 0.95. Llama models correspond to version 3.1 and Turbo means the model is quantized 8bit. Think2SQL-3B and Think2SQL-7B denote the models trained exclusively with RL.
Model Reasoning Simple Medium Challenging Weigthed AVG
Open-source LLMs ($<$ 10B)
DeepSeek-Qwen-1.5B $0.056$ $0.004$ $0.0$ $0.035$
Qwen2.5-Coder-0.5B $0.126$ $0.033$ $0.035$ $0.089$
DeepSeek-Qwen-7B $0.297$ $0.113$ $0.049$ $0.218$
Qwen2.5-Coder-1.5B $0.351$ $0.184$ $0.077$ $0.275$
Llama3-8b $0.436$ $0.260$ $0.133$ $0.355$
Qwen2.5-Coder-3B $0.469$ $0.267$ $0.196$ $0.382$
Qwen2.5-Coder-7B $0.548$ $0.388$ $0.294$ 0.476
Open-source LLMs (10-100B)
DeepSeek-Qwen-32B $0.542$ $0.347$ $0.217$ $0.453$
DeepSeek-Llama-70B $0.552$ $0.371$ $0.203$ $0.465$
QwQ-32B $0.550$ $0.427$ $0.280$ $0.488$
Qwen2.5-Coder-14B $0.610$ $0.456$ $0.364$ $0.541$
Llama-70B-Turbo $0.618$ $0.469$ $0.350$ $0.548$
Qwen2.5-Coder-32B $0.623$ $0.482$ $0.329$ 0.553
Open-source LLMs ( $>$100B)
DeepSeek-R1 $0.588$ $0.440$ $0.294$ $0.518$
Llama-405B-Turbo $0.630$ $0.477$ $0.371$ 0.560
Closed-source LLMs
gpt-4o-mini-2024-07-18 $0.545$ $0.401$ $0.301$ $0.479$
o3-mini-2025-01-31 $0.561$ $0.406$ $0.329$ $0.510$
gpt-4o-2024-08-06 $0.619$ $0.447$ $0.343$ 0.541
Our Models
Think2SQL-3B $0.569$ $0.381$ $0.196$ $0.485$
Think2SQL-7B $0.619$ $0.463$ $0.406$ 0.552
4 Reasoning for Text2SQL
4.1 Experiment Setup
The experiments were designed to evaluate the impact of reasoning on the Text2SQL task. We employed several training strategies, including supervised fine-tuning (SFT), reinforcement learning (RL), and a hybrid approach combining both. Each strategy was designed to assess the influence of reasoning traces on model performance. The research questions we aim to address are:
  • RQ1: Does reasoning improve the performance of Text2SQL models?
  • RQ2: What is the best training strategy to learn reasoning for Text2SQL?
  • RQ3: How do models generalize to unseen databases?
Training and Evaluation datasets. For training, we use two datasets: the original BIRD dataset for RL and the reasoning-augmented BIRD dataset described in Section 3.1 for SFT. The original BIRD dataset is filtered to remove duplicates and erroneous SQL queries, resulting in a cleaned set of $9,007$ examples (after discarding $412$ instances). The reasoning-augmented BIRD dataset contains $1,142$ examples, which are split into $913$ training and $229$ validation samples, following an $80\%/20\%$ ratio.
We evaluate our models on the BIRD development set, which consists of $1,530$ instances (Simple #$\#924$, Medium #$\#461$ and Challenging #$\#143$), as the test set is not publicly available. To assess model robustness and generalization, we also evaluate on the SPIDER dataset [56], a widely recognized benchmark for Text-to-SQL tasks, along with its challenging variants Spider-Syn [14] and Spider-DK [15]. Spider-Syn tests robustness to paraphrased questions by introducing schema-related synonyms, while Spider-DK evaluates the model’s ability to incorporate domain knowledge by modifying both natural language questions and corresponding SQL queries to include implicit relationships or background knowledge not explicitly stated in the schema. For all evaluations, we report Execution Accuracy as the primary metric.
Training setup. To answer the posed research questions, we trained multiple models using different training strategies. Our experiments are based on the Qwen-Coder-2.5 model family [21], focusing specifically on the 3B and 7B variants.
We consider three training approaches: supervised fine-tuning (SFT), reinforcement learning (RL), and a hybrid approach (SFT + RL) that combines both. In the results section, each model is denoted with the respective subscript to indicate the training strategy used.
For SFT, models are trained for 5 epochs using a batch size of 128 and a learning rate of $4\times 10^{-5}$, with the AdamW optimizer [28]. Training is conducted on 4 NVIDIA A100 GPUs, each with 80GB of memory.
For RL, we employ the GRPO algorithm [41] with a batch size of $256$, a learning rate of $1\times 10^{-6}$, and 16 generations per batch, training for 1 epoch. This setup uses 8 NVIDIA H100 GPUs, each also with 80GB of memory.
The hybrid approach (SFT + RL) involves initializing the RL training from a model that has been previously fine-tuned with SFT.
During both training and evaluation, we use the same prompt in Figure 1 and restrict the database schema to include only the tables relevant to the given question. This design choice is deliberate: it helps isolate the reasoning capabilities of the model by removing the confounding influence of schema linking. This allows us to more directly assess the model’s ability to generate SQL from natural language. Our approach aligns with recent work in the Text2SQL domain [7, 3], where SQL generation is treated separately from schema linking.
Model Baselines. To validate our results, we compare a range of open- and closed-source models, with and without reasoning capabilities. Table 1 summarizes the selected models, which vary in size and architecture. All models are evaluated in zero-shot mode using the same training prompt (Figure 1), with a temperature of $0.7$, top_p of $0.95$, and a 30k token generation limit.
We include two main model families: Qwen-Coder-2.5 and LLaMA 3.1 [20]. Qwen-Coder-2.5 allows comparison between reasoning (our) and non-reasoning variants. We also include the recent general-purpose reasoning model from the Qwen family; QwQ [47]. The LLaMA family is used to benchmark DeepSeek-R1 distilled versions against their source models. We also include the 405B LLaMA and the 671B DeepSeek-R1 models.
Among closed-source models, we include o3-mini [35], GPT-4o [34], and its mini variant [33]. Models over 70B parameters are evaluated via Together-AI, and closed-source models via the OpenAI API.
4.2 Main Results
The Table 1 presents a performance comparison of various open-source and proprietary models on the Bird Dev dataset for the Text2SQL task, categorized by model size and reasoning capability. Among open-source models, the Qwen2.5-Coder-32B achieved the highest weighted average accuracy (0.553) in the 10-100B category, while Llama-405B-Turbo led the $>$100B category with a weighted average of 0.560. Proprietary models also performed competitively, with gpt-4o-2024-08-06 achieving a weighted average of 0.541. The Think2SQL models demonstrate strong performance across all evaluation categories, particularly on challenging examples. Think2SQL-7B achieves a weighted average score of 0.552, ranking second overall among all models tested—outperforming both open-source and closed-source models of significantly larger size, including Llama-70B-Turbo and multiple GPT-4 variants. Notably, it achieves the highest score on the Challenging subset (0.406), indicating superior reasoning and generalization abilities. Think2SQL-3B also performs competitively, with a weighted average of 0.485, surpassing all models below 10B parameters and several larger models, such as QwQ-32B and DeepSeek-Llama-70B.
When directly compared to their non-reasoning counterparts of similar size, the Think2SQL models consistently outperform them. Think2SQL-3B exceeds Qwen2.5-Coder-3B in all difficulty categories, with a weighted average of 0.485 ($+0.10$). Likewise, Think2SQL-7B outperforms Qwen2.5-Coder-7B across the board, particularly on moderate and challenging instances-0.388 vs. 0.463 ($+0.07$) and 0.294 vs 0.406 ($+0.11$), respectively. These improvements highlight the effectiveness of incorporating reasoning in training and underscore the competitive edge of our models even when compared to state-of-the-art baselines with similar architectures and parameter counts.
Reasoning capabilities do not always lead to improved performance. For instance, all the distilled DeepSeek models with reasoning perform worse than their non-reasoning counterparts. This suggests that generic reasoning skills are not sufficient to solve the Text2SQL task effectively. Instead, these results highlight the importance of task-specific training: models must be explicitly exposed to structured reasoning within the domain in order to learn how to apply reasoning effectively. Without this targeted supervision, even models equipped with general reasoning abilities may struggle to generalize to complex, domain-specific queries.
Takeaway 1: Think2SQL Incorporating reasoning significantly enhances model performance on the Text2SQL task, particularly for challenging examples. However, general reasoning capabilities alone are insufficient. Task-specific reasoning traces are essential for better performance.
Table 2: Ablation over different training strategies
Model Simple Medium Challenging AVG
Qwen2.5-Coder-3B $0.469$ $0.267$ $0.196$ $0.382$
Qwen2.5-Coder-3B-SFTNO-THINK $0.510$ $0.325$ $0.224$ $0.427$
Qwen2.5-Coder-3B-SFT $0.531$ $0.366$ $0.301$ $0.460$
Qwen2.5-Coder-3B-RLEX $0.569$ $0.381$ $0.273$ $0.485$
Qwen2.5-Coder-3B-SFT-RLEX $0.560$ $0.370$ $0.343$ $0.482$
Qwen2.5-Coder-7B $0.548$ $0.388$ $0.294$ $0.476$
Qwen2.5-Coder-7B-SFT $0.548$ $0.403$ $0.287$ $0.479$
Qwen2.5-Coder-7B-RLEX 0.619 0.463 0.406 0.552
Qwen2.5-Coder-7B-SFT-RLEX $0.590$ $0.422$ $0.343$ $0.516$
4.3 Ablation on different training strategies
Table 2 presents an ablation study comparing different training strategies applied to Qwen2.5-Coder models of two sizes (3B and 7B parameters). Across both model sizes, we observe that applying SFT with reasoning traces consistently improves performance over the base model. Notably, the SFTNO-THINK variant, which lacks reasoning traces, shows a significant drop in performance compared to the full SFT model, particularly on the more challenging examples. For the 3B model, the SFT variant outperforms the SFTNO-THINK of 0.04 points on average, indicating that the reasoning traces are beneficial for the model’s performance.
The model trained only with RL further boosts performance, especially on the more challenging subsets. The combination of SFT and RL (SFT-RLEX) leads to robust and balanced improvements, although in some cases, pure RLEX slightly outperforms it.
For the larger Qwen2.5-Coder-7B model, we see a clear performance gain over its 3B counterpart across all difficulty levels, indicating that model scale remains a strong factor especially for RL when trained with GRPO. The best overall result is obtained with Qwen2.5-Coder-7B-RLEX, achieving the highest average score (0.552) and the strongest performance on Simple and Medium and Challenging examples.
Takeaway 2: RL vs SFT SFT improves over base models, especially for smaller LLMs. RL improves performance across all difficulty levels, particularly in more complex scenarios.
Table 3: Analysis robustness for different datasets.
Model Spider EXSpider-Syn EXSpider-DK EX%
Qwen2.5-Coder-3B $0.725$ $0.632$ $0.602$
Qwen2.5-Coder-3B-SFT $0.770$ $0.718$ $0.634$
Qwen2.5-Coder-3B-RLEX $0.777$ $0.717$ 0.680
Qwen2.5-Coder-3B-SFT-RLEX 0.783 0.760 $0.658$
Qwen2.5-Coder-7B $0.776$ $0.703$ $0.652$
Qwen2.5-Coder-7B-SFT $0.799$ $0.774$ $0.652$
Qwen2.5-Coder-7B-RLEX $0.804$ $0.774$ 0.707
Qwen2.5-Coder-7B-SFT-RLEX 0.824 0.780 $0.673$
4.4 Reasoning robustness on different datasets
Table 3 reports EXof various training strategies across different dataset variants: the original Spider dataset, its synonym-augmented version (Spider-Syn), and the more challenging domain-knowledge variant (Spider-DK). We observe that all training strategies consistently outperform the base Qwen2.5-Coder models for both the 3B and 7B sizes. This confirms the generalizability of our training strategies.
The best overall performance is achieved by the combined SFT-RLEX strategy, which yields the highest accuracy on Spider and Spider-Syn datasets, demonstrating stronger generalization when reasoning traces are incorporated during training. However, on Spider-DK, which demands deeper domain knowledge, pure RL appears to offer an advantage—RLEX outperforms all other strategies on this subset for both model sizes.
Takeaway 3: Generalization with SFT + RL Combining SFT and RL yields the most generalizable models, excelling across diverse datasets. This suggests that integrating reasoning traces with reinforcement learning enables models to better adapt to complex and unseen scenarios.
5 Conclusions, limitations, and future work
This paper investigated the influence of reasoning capabilities on the performance of LLMs for the Text2SQL task. We evaluated different training strategies—Zero-Shot Learning (ZSL) with and without general-purpose reasoning, Supervised Fine-Tuning (SFT) with and without task-specific reasoning traces, Reinforcement Learning (RL) optimizing execution accuracy, and a combined SFT+RL approach—across multiple benchmark datasets.
Our findings answer three research questions. RQ1: while general-purpose reasoning in pretrained LLMs offers limited benefits for complex Text2SQL under ZSL, incorporating task-specific reasoning traces via SFT significantly improves performance, particularly for smaller models. RQ2: RL proved highly effective across all models and datasets, especially for queries demanding multi-hop reasoning - pure RL often yielded the best performance on challenging subsets. RQ3: the combined SFT+RL strategy demonstrated strong generalization across diverse datasets, suggesting it strikes an effective balance between learning general reasoning patterns (via SFT) and optimizing for task-specific correctness (via RL). Our Think2SQL-7B model, trained with RL, achieved performance on par with models exceeding 100 billion parameters on the BIRD dataset, showcasing the power of targeted reasoning reinforcement.
Despite these promising results, this study has limitations. We deliberately isolated the SQL generation process by providing the relevant database schema subset, excluding the challenge of automated schema linking, which is critical in real-world applications. Our RL reward function, while effective, primarily relied on binary execution accuracy, which can be sparse. Furthermore, the SFT training relied on synthetically generated reasoning traces, whose style and quality might influence outcomes. The core experiments focused on the Qwen-Coder-2.5 family, and findings might vary across different model architectures.
Future work should focus on exploring more sophisticated reward modeling for RL, potentially incorporating learned reward models or rewarding partial query correctness to further enhance training efficiency and performance. Evaluating these reasoning-focused training strategies on a wider range of LLM architectures and sizes, as well as on diverse, potentially proprietary or enterprise-specific datasets, would also be valuable. Finally, a deeper qualitative analysis of model failure cases could provide further insights into how different reasoning strategies succeed or falter on specific types of complex queries.